How can I use an identify to find sin(x) and cot(x) if cos(x) = 1/2 and csc(x) < 0?
Answer by Alan3354(69443)

How can I use an identify to find sin(x) and cot(x) if cos(x) = 1/2 and csc(x) < 0?
-------------
sin = sqrt(1 - cos^2)
sine is negative.
-----------
cot = cos/sin